This is a problem from my summer homework that should be really easy, but I'm completely blanking out on how to solve it!
"Christy drove at 50 miles per hour for h hours but arrived g hours late. How fast should Christy have driven to have arrived on time?"
I started some of my own work, but I really had no idea what I was doing, so it was rather pointless, but thanks to anyone who can help!

Found 2 solutions by nerdybill, sudhanshu_kmr:
Answer by nerdybill(7384)

"Christy drove at 50 miles per hour for h hours but arrived g hours late. How fast should Christy have driven to have arrived on time?"
.
Applying the distance formula of d=rt
Distance needed to travel: 50h miles
Time needed to make it in: h-g
.
since:
r = d/t
r = (50h)/(h-g)
.
Therefore, your answer is (relative to h and g):
(50h)/(h-g) mph

Answer by sudhanshu_kmr(1152)


distance = 50 * h = 50h, time = h but g hours late.
actual time require = (h-g) hours
require speed = 50h/(h-g)
difference between speed = 50h/(h-g) - 50 = 50g/(h-g) miles/hour